OpenStudy (anonymous):

Find the geometric mean of the pair of numbers. 36 and 4 A.22 B.17 C.12 D.32

OpenStudy (aravindg):

geometric mean of two numbers a and b is given by \[GM=\sqrt{a \times b}\]

OpenStudy (aravindg):

now can you tell the answer?

OpenStudy (whpalmer4):

Another way of looking at this: \[\frac{4}{x} = \frac{x}{36} \rightarrow x^2 = 4*36 \rightarrow x = ?\]